OpenStudy (anonymous):

I need to be explained this As a salesperson, you are paid $100 per week plus $3 per sale. This week you want your pay to be at least $160. Write an inequality for the number of sales you need to make, and describe the solutions. A) 100 - 3s ≤ 160; s ≤ 20 B) 100 + 3s ≥ 160; s ≥ 18 C) 100 - 3s ≤ 160; s ≤ 18 D) 100 + 3s ≥ 160; s ≥ 20

OpenStudy (anonymous):

Would it be C?

OpenStudy (anonymous):

It would be D

OpenStudy (anonymous):

Ohh okay thank you!

OpenStudy (anonymous):

It see why it make sense

OpenStudy (anonymous):

Your Welcome

OpenStudy (anonymous):

Need Help With Anything Else

OpenStudy (anonymous):

Yes I do actually I have some more question if you don't mind helping :D

OpenStudy (anonymous):

A lion's heart beats 12 times in 16 seconds. How many heartbeats will it have in 60 seconds? A) 3.2 heartbeats B) 36 heartbeats C) 45 heartbeats D) 60 heartbeats

OpenStudy (anonymous):

I would think the answer would be 54 but that's not an answer

OpenStudy (anonymous):

12/16= 3/4 x/60=3/4 Cross multiply. 4x=180 Divide by 4 x=45

OpenStudy (anonymous):

So the answer is 45

OpenStudy (anonymous):

Oh okay so I just didnt use proportions

OpenStudy (anonymous):

Right :)

OpenStudy (anonymous):

I think this one would be c but im not sure...

OpenStudy (anonymous):

Let me see

OpenStudy (anonymous):

Did the file show up?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The Answer Is Indeed C

OpenStudy (anonymous):

The Answer Is

OpenStudy (anonymous):

Subtraction

OpenStudy (anonymous):

Thank you